There are two forms of glucose, designated alpha (a) and beta (b), which are in equilibrium in aqueous solution. The equilibrium constant for the reaction is 1.5 at 30°C.

a-D-glucose(aq) <=> b-D-glucose(aq) K=1.5

(a) If you begin with a fresh 1.0 M solution of a-D-glucose in water, what will be its concentration when equilibrium is reached?

(b) Calculate the percentage of a-glucose and of b-glucose present at equilibrium in aqueous solution at 30°C.
1
Expert's answer
2020-03-24T14:01:11-0400

If the initial concentration of a-gluocse is 1.0 M, in equilibrium solution, the concneration of b-glucose is x while the concentration of a-gluocse is (1-x). As a result, the reaction equilibrium can be described as following:

K = x / (1-x)

where K - the equilibrium constant.

From here:

x = K/ (1+K) = 1.5 / (1 + 1.5) = 0.6 M

1 - x = 0.4 M

(a) Concentration of a-glucose in equilibrium solution is 0.4 M

(b) As 1 M is 100%, the solution consists of 0.4 M × 100% = 40% of a-glucose and 0.6 M × 100% = 60% of b-glucose.


Answer: (a) 0.4 M; (b) 40% and 60%.